Find the domain and range of f(x,y)= 2xy/x^2+y^2


1
Expert's answer
2021-06-10T05:29:57-0400

Let us find the domain and range of "f(x,y)=\frac{2xy}{x^2+y^2}." Since "x^2+y^2=0" implies "(x,y)=(0,0)," we conclude that the domain is "\R^2\setminus\{(0,0)\}." Taking into account that "(x-y)^2\ge0," we conclude that "x^2-2xy+y^2\ge 0," and hence "x^2+y^2\ge 2xy." It follows that "\frac{2xy}{x^2+y^2}\le 1." On the other hand, "(x+y)^2\ge0" implies "\frac{2xy}{x^2+y^2}\ge -1." It follows that the range of "f" is "[-1,1]."